A space vehicle is traveling at 5420 km/h relative to Earth when the exhausted rocket motor is disengaged and sent backward with a speed of 92 km/h relative to the command module. The mass of the motor is four times the mass of the module. What is the speed (in km/h) of the command module relative to Earth just after the separation?
Answer #1
initial velocity vi = 5420 km/h
velocity of the motor vb - va = 92 km/h (relative to the module)
                                 va = vb - 92 km/h
let mass of the command module = m
mass of the motor = 4m
the mass of the vehicle before disengaged = 4m + m = 5m
from law of conservation of momentum ,
      (5m)vi = 4m(va) + m(vb)
      (5m)vi = 4m(vb - 92 km/h) + m(vb)
      (5)vi = 4(vb - 92 km/h) + (vb)
      (5)(5420 km/h) = 4(vb - 92 km/h) + (vb)
     27100 km/h = 4vb +vb - 368 km/h
so ,the speed of the command module
relative to Earth just after the separation is
          vb = 5493.6 km/h
               ˜ 5.493*103 km/h